OpenStudy (anonymous):

bob invests $3700 in an account paying 3% compounded monthly . how much is in the account after 8 months?

OpenStudy (anonymous):

use your eqauation you have all the inputs. principal, rate, and time. i dont remeeber it but if you post it i can give it a shot

OpenStudy (anonymous):

Do you know Compound interest formula?

OpenStudy (anonymous):

is it A= p(1+r/n) ^n*t

OpenStudy (anonymous):

i treid it but i didnt get the right answer

OpenStudy (anonymous):

alrite so , a=3700(1+(.03/n))^n8

OpenStudy (anonymous):

the answer is 3774.65

OpenStudy (anonymous):

Great job at knowing your formula correctly :)

OpenStudy (anonymous):

but i dont kno how to do it

OpenStudy (anonymous):

Compounded monthly -> n = 12

OpenStudy (anonymous):

ok ill try it out

OpenStudy (anonymous):

what do i put for t?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

i still did not get the answer

OpenStudy (anonymous):

Show us how do you calculate it?

OpenStudy (anonymous):

n = 12, n * t = 12 * 8/12 = 8